NEON resumed operations at select sites on May 18, 2020, with required PPE, temperature checks, and other safety precautions in place for staff and surrounding communities. NEON Leadership is continually assessing conditions across the Observatory, resuming operations and reinstating closures where changes in local conditions (i.e., government guidelines, local infection trends, local hospital capacity, and/or local availability of PPE and supplies) dictate.
On March 23, 2020, with the concurrence of the National Science Foundation (NSF), NEON temporarily suspended all activities across NEON that involve in-person or on-site work—including all Domains, headquarters and airborne operations—due to the nationwide risks to health and safety associated with the COVID-19 pandemic. NEON Leadership has approved resumption of operations at offices and sites listed below in Tables 1 and 3 for activities that cannot be conducted from home, following a review of government guidelines, local infection trends, local hospital capacity, local availability of PPE and supplies, and the local operational plan presented by NEON’s local operations managers. Detailed plans have been provided to staff to help ensure their safety and the safety of those around them, including temperature checks and face coverings. The safety of our staff is the top priority, and so end users should anticipate reductions in data collection even at these sites, given the operational restrictions needed to create a safe working environment.
At select Domain support facilities and sites, listed below in Tables 2 and 4, NEON Leadership has authorized resumption of even more limited operations; most NEON sampling and maintenance remain cancelled at these sites. Additional precautions are being taken in these regions to further ensure the safety of our staff in these communities. Locations not listed below are not approved to conduct in-person activities.
As conditions change and governmental requirements indicate, recommendations and decisions may continue to change. NEON Leadership is continually assessing conditions across the Observatory and has reinstated closures where changes in local conditions (i.e., government guidelines, local infection trends, local hospital capacity, and/or local availability of PPE and supplies) indicate increased risk. We will continue to keep you updated on any changes and their impact on Observatory operations. NEON staff remain available to respond to user inquiries.
STATUS (As of 22 February 2021)
|Office Description||Office Location||Approved Restart Date|
|Headquarters - Calibration and Validation Lab, Engineering and Production Lab, Warehouse||Boulder, CO||18 May 2020|
|Hangar||Boulder, CO||18 May 2020|
|Domain 03 Support Facility||Gainesville, FL||22 Feb 2021|
|Domain 04 Support Facility||Guanica, Puerto Rico||9 June 2020|
|Domain 06 Support Facility||Manhattan, KS||22 Feb 2021|
|Domain 09 Support Facility||Jamestown, ND||25 Jan 2021|
|Domain 10/13 Support Facility||Boulder, CO||8 Feb 2021|
|Domain 16 Support Facility||Vancouver, WA||1 Feb 2021|
|Domain 18/19 Support Facility||Fairbanks, AK||1 Feb 2021|
|Domain 20 Support Facility||Hilo, HI||1 June 2020|
|Office Description||Office Location|
|Domain 01 Support Facility||Fitchburg, MA|
|Domain 02 Support Facility||Front Royal, VA|
|Domain 05 Support Facility||Land O'Lakes, WI|
|Domain 07 Support Facility||Oak Ridge, TN|
|Domain 08 Support Facility||Tuscaloosa, AL|
|Domain 12 Support Facility||Bozeman, MT|
|Domain 14 Support Facility||Tucson, AZ|
|Domain 15 Support Facility||Salt Lake City, UT|
|Domain 17 Support Facility||Fresno, CA|
|NEON Domain||Sites with resumed operations||Sites with continuing restrictions||Approved Restart Date|
|D03 - Southeast||OSBS, SUGG, BARC, DSNY, JERC, FLNT||22 Feb 2021|
|D04 - Atlantic Neotropical||GUAN, LAJA, CUPE, GUIL||9 June 2020|
|D06 - Prairie Peninsula||KONZ, KING, KONA, MCDI, UKFS||22 Feb 2021|
|D09 - Northern Plains||WOOD, PRPO, DCFS, PRLA, NOGP||25 Jan 2021|
|D10 - Central Plains||CPER, STER, RMNP, ARIK||8 Feb 2021|
|D13 - Southern Rockies & Colorado Plateau||COMO, WLOU, NIWO||MOAB||8 Feb 2021|
|D16 - Pacific Northwest||MART, MCRA, WREF, ABBY||1 Feb 2021|
|D18 - Tundra||OKSR, TOOL, TOOK||BARR||1 Feb 2021|
|D19 - Taiga||BONA, CARI, DEJU, HEAL||1 Feb 2021|
|D20 - Pacific Tropical||PUUM||8 June 2020|
|D01 - Northeast||HARV, HOPB, BART|
|D02 - Mid-Atlantic||SCBI, POSE, BLAN, LEWI, SERC|
|D05 - Great Lakes||UNDE, CRAM, TREE, STEI, LIRO|
|D07 - Appalachians & Cumberland Plateau||ORNL, WALK, GRSM, LECO, MLBS|
|D08 - Ozarks Complex||BLWA, DELA, LENO, TOMB, MAYF, TALL|
|D12 - Northern Rockies||YELL, BLDE|
|D14 - Desert Southwest||JORN, SRER, SYCA|
|D15 - Great Basin||REDB, ONAQ, (MOAB D13)|
|D17 - Pacific Southwest||SJER, SOAP, BIGC, TEAK, TECR|
Projects that are conducting independent work at NEON sites may be impacted as access to some sites have been temporarily closed per the site host. If you are planning on visiting a NEON site, please double check with the site host or site’s website to determine if the site is accessible. Access to or within NEON infrastructure is not permitted at this time except at the sites listed above (e.g., access to sensor sets, access within fenced areas, or access within NEON sampling plots).
Airborne operations (AOP) will resume in March 2021, with the installation of two payloads into Twin Otter aircraft. Data collection flights will commence according to the schedule below. The safety of each deployment will be evaluated continually, as is true for all of NEON operations, so the proposed schedule below is subject to significant change.
2021 AOP Deployment Schedule - Tentative
|Payload 1||Payload 3|
|Domain||Sched Start Date||Sched End Date||Domain||Sched Start Date||Sched End Date|
|13 (MOAB)||29 Apr||05 May||17 (SJER)||31 Mar||03 Apr|
|7||10 May||29 Jun||14 (SYCA)||05 Apr||11 Apr|
|17 (TEAK, SOAP)||05 Jul||15 Jul||8||21 Apr||08 May|
|16||17 July||27 Jul||15||13 May||23 May|
|2||02 Aug||25 Aug||10 (CPER)||27 May||08 Jun|
|3||27 Aug||03 Oct||9||10 Jun||01 Jul|
|19||07 July||17 Aug|
|14 (SRER)||23 Aug||02 Sep|
|14 (JORN)||04 Sep||07 Sep|
Data Products Status
Observation System (OS) Products
- No new observational data or samples are being collected except at the sites listed above; these data are all dependent on human operators being in the field.
- Samples collected prior to the suspension of activities may continue to be analyzed by external laboratories, if laboratories remain open, resulting in new data becoming available. Samples collected but not yet shipped from field offices remain in storage for later analysis; shipping of samples has resumed where both field offices and the external laboratories have opened.
- Includes DP IDs: DP1.1XXXX, subset of DP1.2XXXX, DP4.00131, DP4.00132
Airborne Observation Products
- Airborne products are being collected according to the tentative schedule above; these data are all dependent on human operators traveling extensively and being in the field.
- Includes DP IDs: DPX.3XXXX
Instrumented System Products
- Most sensors are continuing to collect data and communicate data to the centralized database. There is limited impact to instrumentation currently deployed; however, regular on-site maintenance is suspended.
- Includes DP IDs: DPX.0XXXX (except for DP4.00131, DP4.00132), subset of DP1.2XXXX
- If unplanned power or communications outages occur, NEON Leadership will evaluate each incident on a case-by-case basis to ensure that any efforts by staff to visit the impacted site and restore normal operations can be done safely. Please note that communications outages typically result in delayed delivery of data rather than data loss.
- Sites that are currently experiencing power outages or data loss:
- D11 PRIN (AIS) - Power RESTORED 2021/02/18
- D05 CRAM (AIS) - Power RESTORED 2021/02/14
- D16 MCRA (AIS) - Power down since 2021/02/09
- D11 BLUE (AIS) - Power down since 2020/08/15
- Sites that are currently experiencing communications outages:
- D03 SUGG (AIS) - Communications down since 2021/02/20
- D12 BLDE (AIS) - Communications down since 2021/02/16
- D18 TOOL (TIS) - Communications downs since 2021/02/15
- D11 OAES (TIS) - Communications down since 2021/01/19
- D11 BLUE (AIS) - Water quality data collection has been offline since 2020/08/15 due to power loss stemming from site vandalism. A NEON repair crew is currently deployed to implement upgrades to site infrastructure and restore site power and data collection.
- D08 BLWA (AIS) - Water quality and water surface meteorological data collection has been offline since 2020/01/04 due to mooring failures of the sensor platform (buoy). A new buoy location has been identified and permitted by our site hosts. NEON staff will deploy to the site to reinstall the buoy and resume data collection as COVID-19 and environmental conditions allow. Groundwater and surface water elevation data products continue to be collected.
- D08 TOMB (AIS) - Water quality and water surface meteorological data collection has been offline since 2019/11/02 due to mooring failures of the sensor platform (buoy). NEON is currently exploring sustainable locations for the buoy installation. A team will deploy to reinstall the buoy and resume data collection once a new location has been permitted by our site hosts and COVID-19 conditions allow for our staff to travel safely. Groundwater and surface water elevation data products continue to be collected.
- Primary precipitation sensors (weighing gauge housed within a small double fence intercomparison reference; DP1.00006) that use hazardous chemicals were covered and shut-down at sites at risk of overflow to avoid environmental contamination.
- Currently affected sites: D11 CLBJ, PRIN, BLUE
Wet deposition chemical analysis (DP1.00013) requires regular sample collection from the site and so will not be generated until field operations resume, as captured above.
Eddy Covariance bundled data product (DP4.00200) is expected to continue to be generated during this time. However, NEON staff will remotely monitor the zero-air cylinders that are critical to the safety of the isotopic analyzers. The isotopic analyzers will be shut down if the zero-air cylinders run low. Eddy Covariance Turbulent Exchange and Storage Exchange validations will be shut down automatically as gas cylinders run low, but sampling will be ongoing. Compromised data will be flagged accordingly. No sites have been impacted at this time.
Assignable Asset Status
Battelle staff are continuing to work remotely and the Assignable Asset team will continue to process all current requests and any new requests submitted during this time. PIs with specific questions or concerns about current or planned projects, please contact the Assignable Assets team.